The home appliance department of a large department of store is using the re-order point inventory management system to control the replenishment of a particular model of CD players. The store sells an average of 10 CD players each week. Weekly demand follows a normal distribution with variance 6. Replenishment lead time is 7 weeks. What is the standard deviation of the demand during lead time? Note: Recall from statistics (1) that standard deviation is equal to the square root of variance and (2) that there are only rules for variances but not for standard deviations.  Round your answer to two decimal places.
Answer #1

Variance of demand = 6

So standard deviation of weekly demand (d) = √variance of demand = √6 = 2.45

Lead time (L) = 7 weeks

Standard deviation of demand during lead time = d × √L = 2.45 × √7 = 2.45 × 2.65 = 6.48